Astronomy Report – Mar 25th

Crew 207 Astronomy Report 25 Mar 2019

Crew Astronomer: Tom Baldwin

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested: None

Objects Viewed:

Remote observatory used to image M 65. Image processing not yet complete.

Problems Encountered: None

MUSK OBSERVATORY

Solar Features Observed: None.

Observatory closed, and photos of telescope, Laptop, and Solar Astronomy box sent to Peter.

Problems Encountered: None.

GreenHab Report – March 25th

Crew 207 GreenHab Report – 25 Mar 2019

GreenHab Officer: Diane Rothberg, Julielynn Wong

Environmental control: Heat and cooling w/ambient air 10:00 hr

Shade Cloth (40%)

Average temperatures: RH

Low: 17 deg C 16%

High: 27 deg C 28%

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6 gal

Water in Blue Tank: 136 gallons

Crops watered at 08:00 hours (3 gal) and 17:00 hours (3 gal)

TIME NOTED TEMPERATURE (DEG C) / HUMIDITY (%)

1) 8:00 AM 26 deg C / 25%

2) 10:35 AM 17 deg C / 20%

3) 14:10 PM 27 deg C / 20%

4) 17:00 PM 24 deg C / 20%

Changes to crops:

Narrative: Disposed of a dead pea plant from CRew 206.

Harvest: Basil: 6 gm Parsley 6 gm Sage 1 gm

Support/ Supplies needed: none

Operations Report – Mar 25th

Crew 207 Operations Report 25 Mar 2019

SOL: 9

Name of person filing report: Tom Baldwin

Non-nominal systems:

Generator: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator

Generator (hours run): 12.5 hours

Generator turned on (1930 last PM), SOC (100%)

Generator was turned off at (0700 this AM), SOC (100%)

Diesel Reading – 50 %
Propane Reading – 60 %
Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use
Water (static tank) – 200 gallons (replenished by David)
Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 142 gallons
Water (loft) – Static to Loft Pump used – yes
Water Meter: 142050.1 gallons

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used

Hours:54.9hours

Beginning charge: 100 %

Ending charge: 95 %

Currently charging: Yes (now 100%)

Curiosity rover used: Used

Hours: 84.7 hours

Beginning charge: 100 %

Ending charge: 97 %

Currently charging: Yes (now 100%)

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A
ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where? Not used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts,

Summary of Hab operations: Per recent electrical anomalies, checked SOC of electrical system.

06:00 SOC 100 %, generator on.

07:30 SOC 100%, generator off

10.50 SOC 91%

Summary of Green Hab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Sol Summary Report – Mar 25th

Summary Title : Day #9 (Out of Sim)

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary:

COM continued data collection for the thoracentesis and bike-powered 3D printing studies.

COM + XO continued 3D printing items in Level 1 of the Hab.

ENG + GHO signed out their duties to XO + COM.

ENG + GHO conducted testing of the new drone controller, completed surveys, and submitted qualitative feedback to manufacturer.

XO transported ENG/Astronomer + GHO/Journalist to GJT this PM.

Look ahead plan : Continue 3D printing projects, drone flight mapping, drone controller testing, and data collection for bike-powered 3D printing studies.

Anomalies in work : None

Weather : Sunny

Crew physical status : Fine

EVA : Drone flight testing in PM.

Reports to be filed : Operations, GreenHab, Astronomy

Support Requested : None